### v1

# Neutron Star with Dark Matter Admixture: A Candidate for Bridging the Mass Gap

Abstract
Neutron stars and black holes are the after death remnants of massive stars. However, according to the most recent observations, the neutron stars maximum mass is between $2.0-2.5\ M_{\odot}$ while black holes of less than 5 $M_{\odot}$ rarely has been observed. The region between the most massive neutron star and the least massive black hole is called the mass gap. If indeed its existence is confirmed by future observations, that indicates a gap in our understanding which seeks for explanation. In addition, the existence of compact objects within the mass gap should also be supported with the help of possible new theoretical scenarios. In this letter, we propose a possible explanation for the existence of compact objects within the mass gap region. Specifically, we propose that the mass gap region could be bridged by the existence of a hybrid compact object, composed of hadronic and self interacting - non annihilating fermionic dark matter, considering that the interaction between these two fluids it’s only gravitational. Fundamental questions about how these objects form and how they can be detected are also addressed.
Neutron Stars, Dark Matter, Mass Gap
Astrophysical compact objects come in two varieties: neutron stars (NSs) and black holes (BHs), which are the remnants of massive stars that reach the end of their lives with their cores collapsing in a supernova explosion [1, 2, 3]. NSs are physical nuclear laboratories consist mostly of neutrons, protons and electrons, whereas BHs are too heavy and dense described only by gravity. Both of these entities have uncertain mass limits. In a pioneering idea of Rhoades and Ruffini the optimum upper bound of mass of non-rotating NS was derived using a variations and conditions of standard General-Relativity equation of hydrostatic equilibrium, the Le Chattelier’s principle and the principle of causality [4]. This maximum mass of the equilibrium configuration of a NS cannot be larger than $3.2\ M_{\odot}$[4]. The problem of finding an optimum bound on the mass of non-rotating NS was considered also by other authors (see [5] and references therein) adopting similar assumptions to those of Ref. [4]. However, the majority of realistic nuclear matter equations of state equations predict that NSs maximum mass musdt be between $2.2-2.5\ M_{\odot}$ [1]. In addition, BHs of less than $5\ M_{\odot}$ rarely have been observed. These limits suggest a “mass gap” between the most massive NS and the least massive BH. The existence or not of the mass gap between NS and BH still remain an open problems. While some recent systematic studies support the existence of this gap, there are also recent observations of objects located within this area.
The first event were the mass gap issue started to get interesting was the GW190814 [6] reporting an object with mass $2.6\ M_{\odot}$, later another one, the GW190425 [7] reports a compact binary coalescence with total mass $3.4\ M_{\odot}$ and very recently, the GW230529 [8], came as an addition to the first observations estimating that the one of the components had a mass $3.6\ M_{\odot}$ supporting the possible existence of compact objects close to the lower limit of the mass gap. Also, a number of recent studies (see also Ref.[8]) that have found evidence for compact objects within the mass gap, include observations of non-interacting binary systems [9, 10], radio pulsar surveys [11], and Gravitational Waves (GWs) from compact binary coalescence’s [12, 13, 14, 15]. In any case, an identification whether the component in the gap is a NS or a BH would be remarkable. This distinction is crucial as it will help refine the boundaries of the mass gap, leading to positive advancements in our understanding of the physics of dense nuclear matter[16, 17, 18, 19, 20]. However, it has not yet been conclusively determined whether these components are black holes, neutron stars, or some other type of object. In both cases, one has to resolve this problem. In particular, if this gap really exists then a theoretical explanation for its existence must be given. On the other hand the possible existence of compact objects within this area must also be explained.
Some compact objects observed in GW events have masses inside the gap between known NSs and BHs. The nature of these mass gap objects is unknown, as is the formation of their host binary systems (see also the relevant discussion in Ref.[8]). In fact, there are some theoretical suggestions, including the possibility of exotic compact objects, that could explain these components, such as gravastars [21], boson stars [22], or Planck-scale modifications of BH horizons [23, 24], which could also fall into this gap. Primordial black holes, which may have formed from over-dense regions in the early Universe, have also been proposed as potential candidates to fill the lower mass gap. [25]. In addition to astrophysical and primordial black holes within General Relativity (GR) and exotic compact objects, neutron stars predicted by alternative GR theories could also occupy this mass gap. For instance, in axionic scalar-tensor theory with viable phenomenological equations of state (EoS), such neutron stars could exist within this range. [26].
In the present letter, we propose an alternative explanation involving compact objects within the mass gap region. More specifically, we consider that there could be a stable compact object consisting of NS coexistence with DM within the region in question. This idea, brings to mind the analogy of the DM in Galaxies as an admixture with the baryonic matter, playing a vital role to the rotation curve. Another important question is how DM could become mixed with ordinary matter in a NS. One well studied possibility is through capture as described in the [27, 28, 29, 30, 31, 32, 33, 34, 35]. Also many other possibilities of how celestial bodies like NSs can accumulate DM and interact with neutron matter - nucleons investigated in previous studies [36, 37, 38, 39]. Relevant discussion about the nature, the self-interaction of DM and dark objects are presented in Refs. [44, 41, 45, 46, 43, 47, 48, 42, 40, 50, 51, 49, 52, 53]
In view of the above, a critical and logical question emerges: How can we distinguish these hybrid objects? What sets them apart from a classical NS or a pure DM star? Could the detection of associated gravitational waves (such as through tidal deformability) aid in their identification? The formation and therefore the existence of these objects, could be justified and only astrophysical observations remains to confirm or refute their existence. We also propose some possible ways that may lead to their observational identification.
We consider a compact hybrid object made by two fluids non-self annihilating DM particles admixed with NS matter (including mainly neutrons and a small fraction of protons and electrons). We consider also that the total matter, described by these two fluids, interact each other only gravitationally. In order to predict the bulk properties of the objects consisting of the aforementioned two fluids, one has to solve the coupled Tolman-Oppenheimer-Volkoff (TOV) equations, two for each fluid simultaneously. In the present work, in order to describe the NS matter, we use the EoS developed by Wang et al. [54] in the framework of the Relativistic Brueckner-Hartree-Fock (RBHF) theory. This is an EoS with microscopic origin and its predictions are in very good agreement with both the measured maximum masses and some astrophysical constraints for radii [55]. Now, concerning the DM particles, we consider that they are relativistic fermions, which interact each other through a repulsive force. We consider Yukawa type interaction [40]:
$V(r)=\frac{{\rm g}_{\chi}^{2}(\hbar c)}{4\pi r}\exp\left[-\frac{m_{\phi}c^{2}}{\hbar c}r\right]$ (1)
where ${\rm g}_{\chi}$ is the coupling constant and $m_{\phi}$ is the mediator mass. The role of interaction is decisive for the creation and stability of the mass gap objects. As we mentioned before, the DM is a self-interacting Fermi gas, where the contribution on the energy density of the self-interaction is given by $\frac{y^{2}}{2}(\hbar c)^{3}n_{\chi}^{2}$, where $n_{\chi}$ is the DM density and $y={\rm g}_{\chi}/m_{\phi}c^{2}$ (in units MeV-1), is the interaction strength. The interaction strength $y$ can be constrained by observational limits. These limits imposed on the cross section of the self-interaction [56, 57]. In particular according to [58, 59, 60, 61] it holds $\sigma/m_{\chi}\sim 0.1-10\ {\rm cm}^{2}/{\rm g}$. Moreover, it has been showed that the Born approximation is very accurate in the region $m_{\chi}\sim 1\ {\rm GeV}~{}$[43, 42, 62]. Thus, we found that the interaction parameter $y$ approximately differs across a range spanning two orders of magnitude $\sim(0.001-0.1)({\rm GeV}/m_{\chi}c^{2})^{1/4}({\rm MeV}^{-1})$ [56].
Concerning the stability condition, we employ the method developed by Henriques, Liddle and Moorhouse which used for the study of boson-fermion star [63, 64, 65]. This method has been elaborated and extended through the years for similar studies [66, 68, 69, 67]. According to this method, the stability analysis is carried out by examining the behaviour of baryons and DM particles, yet fixing the total mass $M$ values. In particular, the stability curve is formed with the pair of central values of pressures $\{P_{c}^{\rm NS},P_{c}^{\rm DM}$} exactly in the point where the number of particles reached the minimum and maximum values. This critical curves identify the transition from linear, stable and unstable with respect to perturbations that conserve the mass and the particle number. Hence fulfill the following conditions [67]:
$\left(\frac{\partial N_{b}}{\partial P_{c}^{\rm NS}}\right)_{\rm M=const}=\left(\frac{\partial N_{\chi}}{\partial P_{c}^{\rm NS}}\right)_{\rm M=const}=0$
$\left(\frac{\partial N_{b}}{\partial P_{c}^{\rm DM}}\right)_{\rm M=const}=\left(\frac{\partial N_{\chi}}{\partial P_{c}^{\rm DM}}\right)_{\rm M=const}=0$ (2)
where $N_{b}$ and $N_{\chi}$ are the numbers of baryons and DM particles respectively. The stability analysis can be summarized as follows [67]: Sequences of stable equilibrium configurations that start from purely DM star have the feature that the number of DM particle $N_{\chi}$ first decreases (as a function of the central pressures $P_{c}^{\rm DM}$ or $P_{c}^{\rm NS}$) up to the critical point and the number of baryons $N_{b}$ increases. In general, equilibrium sequences can consist of continuous regions of stability and instability. In this case, therefore, we can have more than one branch of stability.
The number of configurations involving NS and DM mixing in the mass-gap region appears to be infinite. However, in this study, we are limited to only a few cases that can effectively validate our initial hypothesis. The two key parameters of the problem, which remain experimentally undetermined, are the potential mass of DM and the strength of its self-interaction. In fact, the range of these values spans several orders of magnitude for each parameter, and the type of DM - whether fermionic or bosonic - has yet to be determined. However, as we will demonstrate, the overlap range of the two parameters mentioned above, which result in mass-gap configurations, is quite extensive (we also worked with bosonic DM but due to the limitation in length, we present only the fermionic case). A more precise experimental determination of these parameters may also impose limits on the production of such configurations.
Figure 1: Equilibrium configurations of NS and DM admixture, as a function of the central pressures $P_{\rm DM}^{c}$ and $P_{\rm NS}^{c}$, with interaction strength $y=0.05$ MeV-1 and DM mass $m_{\chi}=1500$ MeV. The colored lines correspond to models with the same total mass. The black solid line depicts the boundary between stable and unstable region. The inner figure visualizes more clearly the stable region for very low values of the central pressures.
Figure 2: The figure depicts the radius for the DM component (solid lines), separately with the NS component (dashed lines) and the compactness $\beta$ (dash-doted lines), of the compact object as a function of the central pressure of the NS matter $P_{\rm NS}^{c}$, for a three fixed values of the hybrid compact object.
Figure 3: The dimensional tidal deformability $\Lambda$ of the hybrid compact object as a function of the central pressure of the NS matter $P_{\rm NS}^{c}$, for a three fixed values of the hybrid compact object.
To be more specific, in Fig. 1, we present the equilibrium configurations of a NS with DM admixture and an interaction strength of $y=0.05$ MeV-1 and a DM mass of $m_{\chi}=1500$ MeV. The various colored lines represent models with the same total mass, while the black solid line marks the boundary between the stable and unstable regions. The black dot and the red dot correspond to the purely neutron matter and purely DM star cases (with the corresponding masses). This figure clearly demonstrates that the right combination of NS matter and DM can result in stable compact objects that span the mass-gap range. Another clear conclusion is that, within the aforementioned region, the larger the mass of the star, the higher the proportion of DM it contains. This is expected because, while the equation of state limits the maximum mass of a NS, no such restriction applies to a DM star. Depending on its mass and interactions, a DM star’s mass can grow significantly larger.
Given the importance of their dimensions, particularly for detecting these objects, Fig. 2 illustrates the relationship between the radius of these objects and the central pressure for relatively stable states with defined masses. Specifically, we plot the radius of the DM and NS components separately (with the total radius of the star corresponding to that of the DM component). The radius of the NS core falls within the typical range, while the radius of the DM component is approximately ten times larger. Nevertheless, despite their relatively large halos, these objects remain compact. Specifically, their compactness $\beta=GM/Rc^{2}$, as shown in Fig. 2, is comparable to that of NSs.
However, since the radius of these stars is not easily detectable due to the formation of the DM halo, we focus on the tidal deformability of the star: $\Lambda=\frac{2}{3}k_{2}\left(\frac{c^{2}R}{GM}\right)^{5}$, where $k_{2}$ is the tidal Love number [70]. Thus, in Fig. 3, we plot the dimensional $\Lambda$ for three different masses. Obviously, the values it takes are significantly larger compared to those of NSs. Since $\Lambda$ is a quantity that can be determined from observations, it serves as an important tool for detecting these objects. It is important to note that, for a given mass, the range of $\Lambda$ values can span 2-3 orders of magnitude. This outcome arises from the fact that these objects, in their stable state, can vary significantly in size, particularly in terms of their radius.
Finally, we try to shed light on two critical issues concerning the creation and detection of these objects.When and how could supramassive compact objects made of exotic fermions have formed? A relevant discussion is provided in Ref. [71]. For example, one possibility involves the accretion of DM into NS through the primordial formation of DM clumps surrounded by ordinary matter. Another possibility is the formation of dark compact objects originating from DM perturbations that grow from primordial over-densities.
Finally, another possibility is copious production and capture of DM in the core-collapse supernova of NSs progenitor [72]. In any case, there is no strong theoretical prediction that firmly argues against the creation of these hybrid NS-DM objects. The last proposition leads to the most critical question: how could we locate these objects? The most powerful tool in this case is gravitational lensing, which relies on the space-time distortion caused by these objects. Another possibility is detecting a potential merger event through the well-known detectors of the LIGO-Virgo-KAGRA Collaboration.
This includes two scenarios: a) the merger between two dark compact objects, or b) the merger between one of these dark compact objects and another compact object, such as a NS or a BH. The corresponding gravitational wave signals will carry valuable information about the structure of these objects, providing insights into their composition and internal dynamics [73].
Summarizing this study, we propose an alternative way to explain the possible existence of compact objects in the mass-gap region. We consider that there may be hybrid states of coexisting DM and hadronic matter in the form of stable compact stars in the universe, which, among other things, cover the mass-gap region. The formation of these objects allows for multiple interpretations, and despite its significance, it lies outside the scope of this study.
Acknowledgments — This work was supported by the Czech Science Foundation (GACR Contract No. 21-24281S).